# Towards Interactive and Learnable Cooperative Driving Automation: a Large Language Model-Driven Decision-Making Framework

### v2

@@ Abstract
At present, Connected Autonomous Vehicles (CAVs) have begun to open road testing around the world, but their safety and efficiency performance in complex scenarios is still not satisfactory. Cooperative driving leverages the connectivity ability of CAVs to achieve synergies greater than the sum of their parts, making it a promising approach to improving CAV performance in complex scenarios. However, the lack of interaction and continuous learning ability limits current cooperative driving to single-scenario applications and specific Cooperative Driving Automation (CDA). To address these challenges, this paper proposes CoDrivingLLM, an interactive and learnable LLM-driven cooperative driving framework, to achieve all-scenario and all-CDA. First, since Large Language Models(LLMs) are not adept at handling mathematical calculations, an environment module is introduced to update vehicle positions based on semantic decisions, thus avoiding potential errors from direct LLM control of vehicle positions. Second, based on the four levels of CDA defined by the SAE J3216 standard, we propose a Chain-of-Thought (COT) based reasoning module that includes state perception, intent sharing, negotiation, and decision-making, enhancing the stability of LLMs in multi-step reasoning tasks. Centralized conflict resolution is then managed through a conflict coordinator in the reasoning process. Finally, by introducing a memory module and employing retrieval-augmented generation, CAVs are endowed with the ability to learn from their past experiences. We validate the proposed CoDrivingLLM through ablation experiments on the negotiation module, reasoning with different shots experience, and comparison with other cooperative driving methods. The results demonstrate that our method provides significant safety and efficiency advantages, as well as the ability to interact and learn effectively in complex environments. Our code is available at: https://github.com/FanGShiYuu/CoDrivingLLM.
@@ I Introduction
As autonomous driving technology continues to advance, we are entering an era where both Connected Autonomous Vehicles (CAVs) and Human-Driven Vehicles (HDVs) will coexist. While CAVs are considered to have great potential in improving traffic safety and efficiency, their current performance on open roads is far from satisfactory. According to California’s Department of Motor Vehicles [1], 51% of disengagements were due to CAVs’ decision-making failures. Additionally, the Beijing Autonomous Vehicle Road Test Report revealed that up to 91% of disengagements occurred during interactions with other vehicles, indicating that current autonomous driving technology is not yet adequate for complex interaction scenarios [2].
@@
To ameliorate this issue, a promising approach is to leverage the cooperative driving capabilities of CAVs [3]. The SAE J3216 standard divides Cooperative Driving Automation (CDA) into four levels: Status-sharing, Intent-sharing, Agreement-seeking, and Prescriptive [4]. Federal Highway Administration has conducted the CARMA Program to develop and test CDA. However, the current validation of the program primarily focuses on single CDA under single scenarios, such as cooperative platooning at highway [5]. How to achieve all-scenario and all-CDA cooperative driving still remains to be tackled.
@@
Various approaches have been proposed to address the problem of cooperative driving in different scenarios. The currently popular cooperative driving methods can be broadly categorized into optimization-based, rule-based, and machine learning-based approaches. Optimization-based methods aim to maximize or minimize an objective function to achieve specific goals. Due to the complexity of cooperation involving multiple participants and relationships, optimization-based research often introduces additional simplifications to balance computational time and performance. A common practice in intersection management is to divide the area into multiple zones, with each zone being occupied by only one vehicle at any given time[6, 7]. An alternative approach is to decompose the problem into two layers and optimize the upper and lower layers respectively to simplify the complexity of the problem [8]. However, optimization-based approaches often fail to explicitly consider regulations and social norms during driving, making their decision results difficult to understand. In contrast, the rule-based approach boasts simplicity in form and thus computational efficiency [9]. Nevertheless, despite the rule-based approaches that can be coupled with traffic regulations, pre-set rules often result in poor robustness. On the other hand, machine learning methods, such as deep learning and reinforcement learning, have gained increasing popularity in recent years. These methods have been successfully applied to cooperative driving in environments such as intersections [10, 11, 12], merging zones [13, 14], and highways [15], demonstrating good performance. However, the performance of these learning-based models often degrades significantly when transferred to untrained environments [16, 17]. Overall, although many studies have made numerous attempts at cooperative driving, the significant differences in driving rules and scene characteristics across various scenarios have led to most research focusing on single scenarios. To achieve all-scenario and all-CDA, it is essential to establish a cooperative driving method that possesses both interactive and learnable capabilities.
@@
On the other hand, the rapid development of Transformer models and Large Language Models (LLMs) in recent years has provided a new potential avenue for achieving cooperative decision-making. These models have already shown significant potential in fields such as natural language processing and decision-making for agents. Sha et al. [18] leveraged LLMs’ common sense and high-dimensional information understanding capabilities to handle rare situations, such as a malfunctioning vehicle stalled in the middle of a roundabout. Compared to learning-based and optimization-based single-vehicle decision-making methods, the overall cost was reduced by 18.1% and 16.4%, respectively. Furthermore, to improve the interaction ability of CAV, research has introduced human instructions to establish LLM decision models aligned with human logic and results showed that Human-in-the-loop guidance for LLMs has significantly improved the interpretability of decisions [19, 20]. Building on this, Cui et al. [21] proposed the Talk2Drive framework, the first-of-its-kind multi-scenario field experiment deploying LLMs on a real-world autonomous vehicle. Experimental results showed that Talk2Drive significantly reduced takeover rates in scenarios such as highways, intersections, and parking lots, with at least a 56.6% reduction in takeover rates. In light of the poor performance of current autonomous driving systems in facing long-tail corner cases, Fu et al. [22] proposed an LLM decision-making framework that includes three main functions: reasoning, interpretation, and memorization. This framework enables CAV to dynamically learn through the interaction with the environment, thereby improving its decision-making ability. On this basis, Wen et al. [23] established the Dilu framework, further introducing a reflection module to identify safe and unsafe decisions produced by the reasoning module, subsequently refining them into correct decisions using the knowledge embedded in the LLM, thus further enhancing the learning ability of CAVs. Although LLMs have shown great potential in autonomous driving decision-making tasks, current research is still primarily focused on single-vehicle decision tasks. Designing an efficient, safe, and socially compatible continuous learning framework based on LLMs for multi-vehicle cooperative driving tasks remains a significant area of exploration.
@@
Against this backdrop, given the significance of cooperative driving in addressing the current challenges faced by CAV in complex scenarios and the potential of LLM in enhancing the interaction and learning ability of CAV, this paper proposes an interactive and learnable LLM-driven cooperative driving framework for all-scenario and all-CDA. The contributions of this paper are presented as follows:
Our contributions can be summarized as follows:
@@
  • In the reasoning module, four sub-modules including state-perception, intent-sharing, negotiation, and decision are designed to enable flexible switching between different levels of CDA. Considering the high complexity of multi-vehicle cooperation, a conflict coordinator that infers the suggested passing order for vehicles is designed as the negotiation module. Results show that the introduction of the conflict coordinator significantly improves the vehicle interaction ability.
  • By introducing the memory module with Retrieval Augment Generation (RAG), CAVs are endowed with the ability to continuously learn from historical experiences. By searching for memories most similar to the current scenario and referring to the corresponding decisions, CAVs can avoid repeating past mistakes. Results demonstrated excellent performance across various scenarios and tasks.
@@ II Problem Formulation
With the increasing maturity of autonomous driving technology, manufacturers have shifted from focusing on early-stage technical competition to prioritizing commercial deployment. However, CAVs still exhibit various problems on open roads and even become the culprit of many congestion or accidents. In certain scenarios, the accident rate of CAVs even reached 5.25 times that of human drivers, which has failed to meet people’s expectations for autonomous driving technology, and is gradually eroding people’s trust in CAVs. The communication capabilities of CAVs enable connectivity and mutual assistance. Therefore, leveraging cooperative driving capabilities is a promising way to enhance CAV performance.
Additionally, the cooperative decision-making problem for multiple CAVs can be modeled as a partially observable Markov decision process (POMDP)[24]. We define the POMDP using the tuple $\mathcal{M}_{\mathcal{G}}=(\mathcal{V},S,[\mathcal{O}_{i}],[\mathcal{A}_{i}],\mathcal{P},[r_{i}])$, where $\mathcal{V}$ represents the finite set of all controlled agents (CAVs), and $S$ denotes the state space encompassing all agents. $\mathcal{O}_{i}$ represents the observation space for each agent $i\in\mathcal{V}$, $\mathcal{A}_{i}$ denotes the action space, and $r_{i}$ is the reward associated with CAV $i$. The transition distribution is represented by $\mathcal{P}$.
At any given time, each agent $i$ receives an individual observation $o_{i}:\mathcal{S}\to\mathcal{O}_{i}$ and selects an action $a_{i}\in\mathcal{A}_{i}$ based on a policy $\pi_{i}:\mathcal{O}_{i}\times\mathcal{A}_{i}\to[0,1]$. The agent then transitions to a new state $s_{i}^{\prime}$ with a probability given by $\mathcal{P}(s^{\prime}|s,a):\mathcal{S}\times\mathcal{A}_{1}\times...\times\mathcal{A}_{N}\to\mathcal{S}$.
@@ II-1 Observation Space
Due to the limitations of sensor hardware, a CAV can only detect the status information of surrounding vehicles within a limited distance $\mathcal{L}$. We denote the set of all observable vehicles within the perception range of agent $i$ as $\mathcal{N}_{i}$. The observation matrix for agent $i$, denoted as $\mathcal{O}_{i}$, is a matrix with dimensions $|\mathcal{N}_{i}|\times|\mathcal{F}|$, where $|\mathcal{N}_{i}|$ represents the number of observable vehicles for agent $i$, and $|\mathcal{F}|$ represents the number of features used to describe a vehicle’s state. The feature vector for vehicle $k$ is expressed as
$\mathcal{F}_{k}=[x_{k},y_{k},v^{x}_{k},v^{y}_{k},\cos{\phi_{k}},\sin{\phi_{k}}],$ (1)
where $x_{k}$ and $y_{k}$ are the longitudinal and lateral positions, $v^{x}_{k}$ and $v^{y}_{k}$ are the longitudinal and lateral speeds, and $\cos{\phi_{k}}$ and $\sin{\phi_{k}}$ are the cosine and sine of the vehicle’s heading angle $\phi_{k}$, respectively. The overall observation space of the system is the combined observation of all CAVs, i.e., $\mathcal{O}=\mathcal{O}_{1}\times\mathcal{O}_{2}\times\cdots\times\mathcal{O}_{|\mathcal{V}|}$.
@@ II-2 Action Space
[…]
@@ III Interactive and learnable LLM-driven cooperative driving framework
Leveraging the cooperative driving ability of CAV can significantly improve the safety and efficiency of traffic systems. However, the current research often focuses on single-scenario, single-function cooperation, leading to insufficient interaction capabilities and a lack of continuous learning capabilities. Utilizing the extensive world knowledge and strong reasoning capabilities of LLMs holds promise for addressing these challenges and achieving all-scenario and all-level cooperative automation, thereby fully unleashing the potential of CAVs to transform traffic systems. As a solution, we propose CoDrivingLLM: an interactive and learnable LLM-driven cooperative driving framework for all scenarios and all cooperative driving automation.
@@ III-A Overall Architecture
Fig. 2 illustrates the main modules and their logical relationships within CoDrivingLLM, which consists of three primary modules: the environment module, the reasoning module, and the memory module. First, the environment module updates the current scene information, including the state of all vehicles such as position, speed, etc., based on the actions of both CAVs and HDVs from the previous time step. Next, we designed a centralized-distributed coupled LLM reasoning module. Based on the four levels of CDA defined by the SAE J3216 standard, we integrated four sub-functions into this reasoning module: state sharing, intent sharing, negotiation, and decision. By incorporating the Chain-of-Thought (COT) method, we sequentially connect each sub-function in the reasoning process to enhance the safety and reliability of decision-making. During this process, each CAV performs distributed high-level logical reasoning using LLMs, completing cooperative driving at different levels, and utilizes the conflict coordinator within the framework for centralized conflict resolution, further enhancing safety. Finally, the scenario description, conflict description, and final decisions from the reasoning process are stored in a memory database with a vectorized form. In the subsequent reasoning, the CAV can reference the most similar past memory as experience, enabling the designed CAV to continuously learn and improve its capabilities while driving.
@@ III-B1 Environmental Dynamics Simulation Sub-module
Given that the combination of IDM (Intelligent Driver Model) and MOBIL (Minimizing Overall Braking Induced by Lane changes) is widely used to characterize human driving behavior and has shown good results in various scenarios such as intersections, roundabouts, and merge areas, we utilize IDM and MOBIL to represent the longitudinal and lateral behaviors of HDVs, respectively.
The acceleration of the HDV is given by the IDM as shown below:
$a_{f}(s,v,\Delta v)=a_{\max}\left[1-\left(\frac{v}{v_{\mathrm{d}}}\right)^{\delta}-\left(\frac{s^{*}(v,\Delta v)}{s}\right)^{2}\right]$ (2)
$s^{*}(v,\Delta v)=s_{0}+vT_{g}+\frac{v\Delta v}{2\sqrt{a_{\max}a_{dd}}}$
where $a_{f}(s,v,\Delta v)$ is the acceleration derived from IDM, $v_{\mathrm{d}}$ is the desired velocity, $\delta$ is the acceleration exponent, $\Delta v$ and $s$ are the velocity difference and distance between the front vehicle and the subject vehicle, $s^{*}(v,\Delta v)$ is the expected distance, $s_{0}$ is the minimum stopping distance, $T_{g}$ is the desired time gap, $a_{\max}$ is the maximum acceleration, and $a_{dd}$ is the desired deceleration.
Furthermore, MOBIL achieves safe and efficient traffic flow by minimizing the overall braking caused by lane change, which mainly includes two parts: lane change incentive and safety inspection. The lane change incentive evaluates the change in the acceleration of the lane change on itself and surrounding vehicles to determine whether a lane change is warranted.
$\underbrace{\tilde{a}_{c}-a_{c}}_{\text{ego-vehicle}}+p(\underbrace{\tilde{a}_{n}-a_{n}}_{\text{new follower}}+\underbrace{\tilde{a}_{o}-a_{o}}_{\text{old follower}})\geq\Delta a_{\text{th}},$ (3)
where $\tilde{a}$ and $a$ are the acceleration of the vehicles before and after the lane change, respectively, $p$ is a politeness coefficient that determines the degree to which the vehicle pays attention to the acceleration changes of the surrounding vehicles when making lane change decisions, and $a_{\text{th}}$ the acceleration gain required to trigger a lane change.
At the same time, in order to ensure the safety of lane change, the MOBIL model also carries out safety inspection to ensure that lane change will not cause a sudden brake on the rear vehicle of the expected lane, that is:
$\tilde{a}_{n}\geq-b_{safe},$ (4)
where $b_{safe}$ is the maximum braking imposed on a vehicle during a cut-in.

### v3

@@ Abstract
Connected Autonomous Vehicles (CAVs) are being tested globally, but their performance in complex scenarios remains suboptimal. While cooperative driving improves CAV performance by leveraging vehicle collaboration, its lack of interaction and continuous learning limits current applications to single scenarios and specific Cooperative Driving Automation (CDA). To address these issues, this paper proposes CoDrivingLLM, an interactive and learnable LLM-driven cooperative driving framework for all-scenario and all-CDA applications. First, an environment module updates vehicle positions based on semantic decisions, mitigating errors from LLM-controlled positioning. Second, leveraging the four CDA levels defined in SAE J3216, a centralized-distributed coupled architecture reasoning module is developed to ensure safe and efficient cooperation through centralized negotiation and distributed decision. Finally, by introducing a memory module that employs Retrieval Augmented Generation (RAG), CAVs are endowed with the ability to learn from their past experiences to avoid repeating mistakes. Through ablation studies and comparisons with other cooperative driving methods, the results demonstrate that the proposed CoDrivingLLM significantly enhances safety, efficiency, and adaptability across various scenarios. Our code is available at: https://github.com/FanGShiYuu/CoDrivingLLM.
@@ I Introduction
As autonomous driving technology continues to advance, we are entering an era where both Connected Autonomous Vehicles (CAVs) and Human-Driven Vehicles (HDVs) will coexist. While CAVs are considered to have great potential in improving traffic safety and efficiency, their current performance on open roads is far from satisfactory. According to California’s Department of Motor Vehicles [1], 51% of disengagements were due to CAVs’ decision-making failures. Additionally, the Beijing Autonomous Vehicle Road Test Report revealed that up to 91% of disengagements occurred during interactions with other vehicles, indicating that current autonomous driving technology is not yet adequate for complex interaction scenarios [2, 3].
@@
To ameliorate this issue, a promising approach is to leverage the cooperative driving capabilities of CAVs [4]. The SAE J3216 standard divides Cooperative Driving Automation (CDA) into four levels: Status-sharing, Intent-sharing, Agreement-seeking, and Prescriptive [5]. Federal Highway Administration has conducted the CARMA Program to develop and test CDA. The SARTRE project, led by Volvo, has turned cooperative platooning into a reality. However, the current validation of the program primarily focuses on single CDA under single scenarios [6]. How to achieve all-scenario and all-CDA cooperative driving still remains to be tackled.
@@
Various approaches have been proposed to address cooperative driving, typically categorized into optimization-based, rule-based, and machine learning-based methods. Optimization-based methods aim to optimize an objective function, often simplifying the problem to balance performance and computational efficiency. Common strategies include dividing the area into zones, each occupied by one vehicle [7, 8], or decomposing the problem into two layers [9, 10]. However, these methods often neglect driving regulations and social norms, making their decisions harder to interpret. In contrast, the rule-based approach boasts simplicity in form and thus computational efficiency [11, 12]. Chen et al. [13] conducted a comparison between the improved Depth-First Search Tree (iDFST) method in a large-scale intersection where over 30 vehicles were involved. Nevertheless, despite the rule-based approaches that can be coupled with traffic regulations and is suitable to large-scale systems, pre-set rules often result in poor robustness [14]. On the other hand, machine learning methods, such as deep learning and reinforcement learning, have gained increasing popularity in recent years. These methods have been successfully applied to cooperative driving in environments such as intersections [15, 16], merging zones [17, 18], and highways [19], demonstrating good performance. Bae et al. [20] combined RNN with MPC for cooperative driving in high-density scenarios, achieving strong real-time performance. Zhou et al. [15] utilized Multi-Agent Deep Deterministic Policy Gradient (MADDPG) to model the cooperation among agents and results showed that MADDPG outperforms IDM in fuel and efficiency by 7.4% and 5.3%. However, the performance of these learning-based models often degrades significantly when transferred to untrained environments [21, 22, 23]. Overall, while many studies have explored cooperative driving, the varying scene characteristics have led most research to focus on a single scenario. To achieve all-scenario and all-CDA, it is essential to establish a method that possesses both interactive and learnable capabilities.
@@
Additionally, recent advancements in transformer models and Large Language Models (LLMs) have opened new possibilities for cooperative decision-making [24, 25]. These models have demonstrated strong potential in understanding human behavior and interacting with humans. [26, 27, 28]. Sha et al. [29] utilized LLMs’ understanding of common sense and complex scenarios, such as a malfunctioning vehicle in a roundabout, reducing overall costs by 18.1% and 16.4% compared to single-vehicle decision-making methods. Furthermore, to improve the interaction ability of CAV, research has introduced human instructions to establish LLM decision models aligned with human logic and results showed that Human-in-the-loop guidance for LLMs has significantly improved the interpretability of decisions [30, 31]. Building on this, Cui et al. [32] introduced the Talk2Drive framework, a real-world deployment of LLMs, which reduced takeover rates by 56.6% in various scenarios. To enable the LLM to learn from mistakes, Wen et al. [33] introduced a reflection module to refine decisions using LLM knowledge, improving CAV learning. Hu et al. [34] introduced a communication module into LLM, which improved the success rate from 45% to 75%, highlighting the critical role of communication in enhancing CAV efficiency and safety in complex environments. Despite these advancements, most research still focuses on single-vehicle decision-making. Therefore, designing an efficient, safe, and socially compatible continuous learning framework based on LLMs for multi-vehicle cooperative driving tasks remains a significant area of exploration.
@@
Against this backdrop, given the significance of cooperative driving in addressing the current challenges faced by CAV in complex scenarios and the potential of LLM in enhancing the interaction and learning ability of CAV, this paper proposes an interactive and learnable LLM-driven cooperative driving framework for all-scenario and all-CDA. Our contributions can be summarized as follows:
@@
  • In the reasoning module, four sub-modules including state-sharing, intent-sharing, negotiation, and decision are designed to enable flexible switching between different levels of CDA. Considering the high complexity of multi-vehicle cooperation, a conflict coordinator that infers the suggested passing order for vehicles is designed as the negotiation module. Results show that the introduction of the conflict coordinator significantly improves the vehicle interaction ability.
  • By introducing the memory module with Retrieval Augment Generation (RAG), CAVs are endowed with the ability to continuously learn from historical experiences. By searching for memories most similar to the current scenario and referring to the corresponding decisions, CAVs can avoid repeating past mistakes. Results demonstrated that with an increasing number of historical interactions, the success rate of CoDrivingLLM steadily improved across different scenarios.
@@ II Problem Formulation
With the increasing maturity of autonomous driving technology, manufacturers have shifted from focusing on early-stage technical competition to prioritizing commercial deployment. However, CAVs still exhibit various problems on open roads and even become the culprit of many congestion or accidents. In certain scenarios, the accident rate of CAVs even reached 5.25 times that of human drivers [35], which has failed to meet people’s expectations for autonomous driving technology, and is gradually eroding people’s trust in CAVs. The communication capabilities of CAVs enable connectivity and mutual assistance. Therefore, leveraging cooperative driving capabilities is a promising way to enhance CAV performance.
Additionally, the cooperative decision-making problem for CAVs can be modeled as a Partially Observable Markov Decision Process (POMDP) [36]. We define the POMDP using the tuple $\mathcal{M}_{\mathcal{G}}=(\mathcal{V},\mathcal{S},[\mathcal{O}_{i}],[\mathcal{A}_{i}],\mathcal{P})$, where:
  • $\mathcal{V}$ represents the finite set of all controlled agents (CAVs),
  • $\mathcal{S}$ denotes the state space, encompassing the states of all agents and the environment,
  • $\mathcal{O}_{i}$ represents the observation space for each agent $i\in\mathcal{V}$,
  • $\mathcal{A}_{i}$ denotes the action space for agent $i$,
  • $\mathcal{P}$ represents the transition function, capturing the probability of moving from one state to another.
At any given time step, each agent $i$ receives an individual observation $o_{i}:\mathcal{S}\to\mathcal{O}_{i}$ and selects an action $a_{i}\in\mathcal{A}_{i}$ based on a policy $\pi_{i}:\mathcal{O}_{i}\times\mathcal{A}_{i}\to[0,1]$. The agent then transitions to a new state $s_{i}^{\prime}$ with a probability given by the state transition function $\mathcal{P}(s^{\prime}|s,a):\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{N}\to\mathcal{S}$, where $N$ is the total number of agents.
@@ II-1 Observation Space
Due to the limitations of sensor hardware, a CAV can only detect the status information of surrounding vehicles within a limited distance $\mathcal{L}$. We denote the set of all observable vehicles within the perception range of agent $i$ as $\mathcal{N}_{i}$. The observation space for agent $i$, denoted as $\mathcal{O}_{i}$, is a matrix with dimensions $|\mathcal{N}_{i}|\times|\mathcal{F}|$, where $|\mathcal{N}_{i}|$ represents the number of observable vehicles for agent $i$, and $|\mathcal{F}|$ represents the number of features used to describe a vehicle’s state. The feature vector for vehicle $i$ is expressed as:
$\mathcal{F}_{i}=[x_{i},y_{i},v^{x}_{i},v^{y}_{i},lane_{i},int_{i}]$ (1)
where $x_{i}$ and $y_{i}$ are the longitudinal and lateral positions, $v^{x}_{i}$ and $v^{y}_{i}$ are the longitudinal and lateral speeds, $lane_{i}$ represents the lane information of vehicle $i$, and $int_{i}$ is the driving intention of vehicle $i$. $lane_{i}$ and $int_{i}$ will be described in detail in the next section.
The overall observation space of the system is the combined observation of all CAVs, i.e.,
$\mathcal{O}=\mathcal{O}_{1}\times\mathcal{O}_{2}\times\cdots\times\mathcal{O}_{|\mathcal{V}|}.$
@@ II-2 Action Space
Given that the strength of LLMs lies in their reasoning capabilities based on world knowledge rather than numerical computation, we design the decision actions of CAVs as discrete semantic decisions rather than direct vehicle control actions. The action space $\mathcal{A}_{i}$ for agent $i$ is defined as a set of high-level control decisions, including $\{\textit{SLOWER},\textit{IDLE},\textit{FASTER},\textit{LANE LEFT},\textit{LANE RIGHT}\}$. Once a high-level decision is selected, lower-level controllers generate the corresponding steering and throttle control signals to manage the CAVs’ movements. The overall action space is the combination of actions from all CAVs, i.e.,
$\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\times\cdots\times\mathcal{A}_{|\mathcal{V}|}.$
II-3 Transition Dynamics
The transition dynamics of the system, encapsulated in the transition function $\mathcal{P}$, are modeled by the kinematics of the vehicles. The motion of each CAV is influenced by both the decision actions of the agent and the behaviors of other surrounding vehicles, particularly HDVs in mixed traffic environments. The state update for each CAV is based on a combination of high-level semantic decisions and low-level control signals generated by the vehicle’s control system.
@@ III Interactive and learnable LLM-driven cooperative driving framework
Leveraging the cooperative driving ability of CAV can significantly improve the safety and efficiency of traffic systems. However, the current research often focuses on single-scenario, single-function cooperation, leading to insufficient interaction capabilities and a lack of continuous learning capabilities. Utilizing the extensive world knowledge and strong reasoning capabilities of LLMs holds promise for addressing these challenges and achieving all-scenario and all-CDA, thereby fully unleashing the potential of CAVs to transform traffic systems. As a solution, we propose CoDrivingLLM: an interactive and learnable LLM-driven cooperative driving framework for all-scenarios and all-CDA.
@@ III-A Overall Architecture
Fig. 2 illustrates the main modules and their logical relationships within CoDrivingLLM, which consists of three primary modules: the environment module, the reasoning module, and the memory module. First, the environment module updates the current scene information, including the states of all vehicles such as position, speed, etc., based on the actions of both CAVs and HDVs from the previous time step. Next, we design a centralized-distributed coupled LLM reasoning module. Based on the four levels of CDA defined by the SAE J3216 standard, we integrate four sub-functions into this reasoning module: state sharing, intent sharing, negotiation, and decision. By incorporating the Chain-of-Thought (COT) method, we sequentially connect each sub-function in the reasoning process to enhance the safety and reliability of decision-making. During this process, each CAV uses LLM for distributed high-level logical reasoning and completes cooperative driving at various levels. Then, a conflict coordinator within the framework handles centralized conflict resolution, further improving safety. Finally, the scenario description, conflict description, and final decisions from the reasoning process are stored in a memory database with a vectorized form. In the subsequent reasoning, the CAV can reference the most similar past memory as experience, enabling the designed CAV to continuously learn and improve its capabilities while driving.
@@ III-B1 Environmental Dynamics Simulation Sub-module
Given that the combination of IDM (Intelligent Driver Model) and MOBIL (Minimizing Overall Braking Induced by Lane changes) is widely used to characterize human driving behavior and has shown good results in various scenarios such as intersections, roundabouts, and merge areas, we utilize IDM and MOBIL to represent the longitudinal and lateral behaviors of HDVs, respectively. Since IDM and MOBIL are not the methodological contributions of this study, the detailed descriptions of these models can be found in the appendix here  [^1] https://github.com/FanGShiYuu/CoDrivingLLM/blob/master/highway_env/HDV.md